OpenStudy (anonymous):

Find the value of the annuity at the end of the indicated time period. The amount, frequency of deposits, annual interest rate, and time period are given. Do not round until the final answer. Then, round to the nearest cent. Amount: $225; quarterly; 9%; 8 years.

OpenStudy (amistre64):

the net present value or the future value?

OpenStudy (anonymous):

future value

OpenStudy (anonymous):

A $10,381.03 B $10,869.74 C $11,169.84 D $16,469.85

OpenStudy (amistre64):

the future value is just a normal compounding interest setup

OpenStudy (amistre64):

\[B_t=B_o(1+\frac{r}{c})^{ct}+P\frac{1-(1+\frac{r}{c})^{ct}}{1-(1+\frac{r}{c})}\]

OpenStudy (amistre64):

since Bo is 0, that amounts to the Payments only

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (amistre64):

youre welcome